i have a square inside a circle how do i find its area when they only give me the base and height of the sqaur
Answer by lwsshak3(11628)

i have a square inside a circle how do i find its area when they only give me the base and height of the square.
The diagonal of the square is equal to the diameter of the circle.
if x=one side of square
diagonal=diameter of circle=√(2x^2) (pythagorean theorem)
radius=√(2x^2)/2
Area of circle=πr^2=π2x^2/4=πx^2/2